# GenBench: A Benchmarking Suite for Systematic Evaluation of Genomic Foundation Models

### v1

@@ Abstract
The Genomic Foundation Model (GFM) paradigm is expected to facilitate the extraction of generalizable representations from massive genomic data, thereby enabling their application across a spectrum of downstream applications. Despite advancements, a lack of evaluation framework makes it difficult to ensure equitable assessment due to experimental settings, model intricacy, benchmark datasets, and reproducibility challenges. In the absence of standardization, comparative analyses risk becoming biased and unreliable. To surmount this impasse, we introduce GenBench, a comprehensive benchmarking suite specifically tailored for evaluating the efficacy of Genomic Foundation Models. GenBench offers a modular and expandable framework that encapsulates a variety of state-of-the-art methodologies. Through systematic evaluations of datasets spanning diverse biological domains with a particular emphasis on both short-range and long-range genomic tasks, firstly including the three most important DNA tasks covering Coding Region, Non-Coding Region, Genome Structure, etc. Moreover, We provide a nuanced analysis of the interplay between model architecture and dataset characteristics on task-specific performance. Our findings reveal an interesting observation: independent of the number of parameters, the discernible difference in preference between the attention-based and convolution-based models on short- and long-range tasks may provide insights into the future design of GFM.
@@ 1 Introduction
Despite the importance of modeling Genomics foundations and the advancement of different training methods, there is still a noticeable absence of a thorough benchmark in this area that encompasses a range of practical application scenarios and different foundational model structures. The current benchmark either restricts its scope to short distances or oversimplifies the challenge by focusing solely on the classification task. Moreover, with the influx of long sequence models [4, 5, 6, 7], a systematic approach to evaluating up-to-date GFMs and inspiring subsequent development is also sorely lacking. Based on the current state of research, we thus summarise the three key problems in current GFMs: (1) Incomplete evaluation: Long sequence processing is crucial for modeling genetic data. Current tests of these models for long-sequence gene tasks are incomplete. (2) Chaotic training strategies: The variety of tokenizations and pre-training methods lack a fair platform to compare and select the most appropriate training method for DNA data. (3) Snowy model design: Do the various Attention-based and recent state-space models each have strengths in DNA? We need an experience that inspires subsequent designs.
Figure 1: The three important parts in the organization of eukaryotic genomic DNA: Coding Region, Non-coding Region, and Genome Architecture. The small proportion of coding regions (1.5%) that code for proteins, in contrast to the vast non-coding regions (98.5%) involved in gene regulation. As a result, DNA is wrapped around histone proteins and densely packed as a chromosome, also named Genome Architecture. Notably, long-range interactions in DNA are very important in life processes.
@@ 3.2 Support Tasks
Long-Range Tasks. Long-range tasks are defined as tasks with input lengths longer than 1000. Achieving state-of-the-art performance on benchmarks involving long sequences, such as the Long Range Arena (LRA) [19], is possible. However, longer context lengths also introduce higher complexity. For instance, attention-based models exhibit quadratic complexity concerning input length [20]. In long-range tasks, we include site annotation [14], species classification [6], promoter prediction [2], chromatin profiling [21], and genomic structure prediction [22]. These tasks cover a diverse range from multi-class classification to regression.
@@ 3.4 Codebase Structure
Dataset HyenaDNA($\uparrow$) DNABERT($\uparrow$) DNABERT2($\uparrow$) GENA-LM($\uparrow$) NT($\uparrow$) Caduceus($\uparrow$) CNN($\uparrow$) Mouse Enhancers 0.7934 0.8099 0.8182 0.8297 0.8512 0.8163 0.7008 Coding vs Intergenomic 0.9097 0.9364 0.9358 0.9324 0.9576 0.9372 0.8844 Human vs Worm 0.9624 0.9584 0.9739 0.9698 0.9751 0.9557 0.9408 Human Enhancers Cohn 0.7296 0.7023 0.7587 0.7563 0.7612 0.7376 0.7080 Human Enhancers Ensembl 0.9033 0.8919 0.9075 0.9107 0.9244 0.8448 0.7637 Human Ensembl Regulatory 0.8462 0.938 0.8832 0.8810 0.9403 0.7367 0.8616 Human Nontata Promoters 0.9445 0.8713 0.9524 0.9660 0.9295 0.8885 0.8564 Human OCR Ensembl 0.7914 0.7496 0.7582 0.7898 0.8042 0.8176 0.6947 Human Core Promoter Detection 0.8440 0.8491 0.8257 0.8140 0.8541 0.8505 0.8003 Human Transcription Factor Prediction 0.6976 0.7840 0.8218 0.8240 0.8262 0.6928 0.6672 Human Promoter Detection 0.7295 0.8393 0.8993 0.9001 0.939 0.7322 0.6875 Human Splice Site Detection 0.5660 0.8721 0.8813 0.9178 0.9481 0.5674 0.5666 Mouse Transcription Factor Prediction 0.6535 0.7393 0.8269 0.8265 0.8502 0.6519 0.6081 Yeast Epigenetic Marks Prediction 0.6301 0.7203 0.8022 0.7829 0.7845 0.6378 0.6071 Virus Covid Variant Classification 0.3770 0.5990 0.7195 0.7033 0.6939 0.3794 0.1974
@@ 4.1 Short Range Tasks
Convolution-based models consistently show lower performance compared to attention-based models, particularly on challenging tasks. In binary classification scenarios, HyenaDNA demonstrates a decrease in accuracy of 0.053 compared to the attention-based model. This discrepancy widens to 0.239 in multi-class classification tasks. Similarly, Caduceus exhibits a 0.053 accuracy gap in binary tasks and a larger margin of 0.274 in multi-class classification assignments. In comparison to the pre-trained model, CNN achieved significantly lower accuracy, with a gap up to 0.387.
@@ 4.3 Effect of Length
We analyzed the impact of sequence length on performance in long-range tasks. We utilized input sequences of lengths 512, 1000, 2000, and 3000 base pairs(bp) for Hyena-DNA, Nucleotide Transformer, DNABERT-2, and GENA-LM in species and promoter prediction tasks. To investigate the feasibility of convolution-based models, we tested a 30,000 base pair input sequence for Hyena-DNA specifically in the species prediction task. The results are presented in top right and left of Figure 2. It is evident that performance improvement is consistent with increasing sequence length across all models. While Hyena-DNA may lag behind other attention-based models at shorter context lengths, it achieve the best performance in longer contexts, highlighting the advantages of longer context lengths. However, in datasets like promoter prediction, where the input length is capped at 3000 bp, this limitation poses a challenge for future research.
@@ 4.4 Gene Clustering
In this section, we examine the fine-tuned embedding models HyenaDNA, DNABERT2, GENA-LM, Caduceus, and Nucleotide Transformer. These models encode gene sequences from various species. We extract the representations from the final hidden layer and employ t-distributed Stochastic Neighbor Embedding (t-SNE) [26] for visualization. The visualization in Figure 3 illustrates clear clusters visually and provides quantitative insights. Specifically, for the NT, which exhibits the highest accuracy, the embeddings of distinct species are visibly separated. In contrast, for HyenaDNA, which has the lowest accuracy, there is less differentiation among the embeddings of different species.
@@ 5 Conclusion and Dicussion
Acknowledgements
This work was supported by Ministry of Science and Technology of the People’s Republic of China (No. 2021YFA1301603), National Natural Science Foundation of China Project (No. U21A20427), Project (No. WU2022A009) from the Center of Synthetic Biology and Integrated Bioengineering of Westlake University and Project (No. WU2023C019) from the Westlake University Industries of the Future Research Funding. This work was done when Jiahui Li interned at Westlake University. We thank the AI Station of Westlake University for the support of GPUs.
@@ Appendix A Codebase Overview
In this section, we present a comprehensive overview of the codebase structure of GenBench. The codebase is organized into three abstracted layers, namely the core layer, algorithm layer, and user interface layer, arranged from the bottom to the top, as illustrated in Figure 5
@@
Core Layer The core layer of GenBench includes key elements like data loaders for supported datasets, fundamental modules for supported models, and metrics for evaluation. Data loaders provide a standardized way for loading and preprocessing data. The modules contain essential unit implementations of supported models. Metrics offer a consistent method for evaluating results. This core layer sets the groundwork for upper layers to ensure adaptable usage.
@@ Appendix B Implementation Details
The table presented in Table 9 outlines the hyperparameters utilized in the various models supported across different datasets. Each model’s hyperparameters consist of layers, width of the hidden dimension, parameter size, learning rate, embed dropout, residual dropout, optimizer, optimizer momentum, training epochs, batch size, LR scheduler, and reverse complement augmentation. It is important to note that sequence length is task-dependent and not directly related to the model.
B.1 Data Description
Moreover, we have integrated genomic structure prediction as presented in [17]. This prediction specifically examines how structural variants impact genome organization at various scales. In order to enhance practicality, we have limited the length to 6000-bp. Additionally, we have included Species Classification from [6], which has heightened the complexity of classification by encompassing a larger number of species.
We incorporated the task of Bulk RNA expression to evaluate the model’s performance within a lengthy context [28]. Bulk RNA-sequencing is a biological assay that gauges the average gene expression from a group of cells within a specific tissue. In this particular task, the input is the sequence from the human reference genome that is centered around the Transcription Start Site (TSS). The resulting output is a unified vector consisting of continuous values that indicate the bulk RNA levels of a gene across 218 distinct tissue types.
B.2 Model Description

### v2

@@ Abstract
The Genomic Foundation Model (GFM) paradigm is expected to facilitate the extraction of generalizable representations from massive genomic data, thereby enabling their application across a spectrum of downstream applications. Despite advancements, a lack of evaluation framework makes it difficult to ensure equitable assessment due to experimental settings, model intricacy, benchmark datasets, and reproducibility challenges. In the absence of standardization, comparative analyses risk becoming biased and unreliable. To surmount this impasse, we introduce GenBench, a comprehensive benchmarking suite specifically tailored for evaluating the efficacy of Genomic Foundation Models. GenBench offers a modular and expandable framework that encapsulates a variety of state-of-the-art methodologies. Through systematic evaluations of datasets spanning diverse biological domains with a particular emphasis on both short-range and long-range genomic tasks, firstly including the three most important DNA tasks covering Coding Region, Non-Coding Region, Genome Structure, etc. Moreover, We provide a nuanced analysis of the interplay between model architecture and dataset characteristics on task-specific performance. Our findings reveal an interesting observation: independent of the number of parameters, the discernible difference in preference between the attention-based and convolution-based models on short- and long-range tasks may provide insights into the future design of GFM. We made the GenBench codebase and associated models publicly accessible here.
@@ 1 Introduction
Despite the importance of modeling Genomics foundations and the advancement of different training methods, there is still a noticeable absence of a thorough benchmark in this area that encompasses a range of practical application scenarios and different foundational model structures. The current benchmark either restricts its scope to short distances or oversimplifies the challenge by focusing solely on the classification task [1, 4, 2]. Moreover, with the influx of long sequence models [5, 6, 7, 8], a systematic approach to evaluating up-to-date GFMs and inspiring subsequent development is also sorely lacking. Based on the current state of research, we thus summarise the three key problems in current GFMs: (1) Incomplete evaluation: Long sequence processing is crucial for modeling genetic data. Current tests of these models for long-sequence gene tasks are incomplete. (2) Chaotic training strategies: The variety of tokenizations and pre-training methods lack a fair platform to compare and select the most appropriate training method for DNA data. (3) Snowy model design: Do the various Attention-based and recent state-space models each have strengths in DNA? We need an experience that inspires subsequent designs.
Figure 1: The image illustrates the intricate organization of eukaryotic genomic DNA, highlighting three critical components: the Coding Region, the Non-coding Region, and Genome Architecture. The Coding Region, comprising only about 1.5% of the genome, is responsible for coding proteins, essential for cellular functions. In contrast, the vast Non-coding Region, making up 98.5% of the genome, plays crucial roles in gene regulation and genome stability, containing regulatory elements like promoters, enhancers, and silencers, as well as structural components. Genome Architecture is depicted by the DNA being tightly packed around histone proteins into chromosomes, ensuring efficient storage and regulation of genetic information. Notably, long-range interactions within the DNA are essential for life processes, facilitating the regulation of gene expression by bringing distant regulatory elements into proximity with target genes.
@@ 3.2 Support Tasks
Long-Range Tasks. Long-range tasks are defined as tasks with input lengths longer than 1000. Achieving state-of-the-art performance on benchmarks involving long sequences, such as the Long Range Arena (LRA) [20], is possible. However, longer context lengths also introduce higher complexity. For instance, attention-based models exhibit quadratic complexity concerning input length [21]. In long-range tasks, we include site annotation [15], species classification [7], promoter prediction [2], chromatin profiling [22], and genomic structure prediction [23].
3.3 Data Description
Moreover, we have integrated genomic structure prediction as presented in [18]. This prediction specifically examines how structural variants impact genome organization at various scales. In order to enhance practicality, we have limited the length to 6000-bp. Additionally, we have included Species Classification from [7], which has heightened the complexity of classification by encompassing a larger number of species. We incorporated the task of Bulk RNA expression to evaluate the model’s performance within a lengthy context [25]. Bulk RNA-sequencing is a biological assay that gauges the average gene expression from a group of cells within a specific tissue. In this particular task, the input is the sequence from the human reference genome that is centered around the Transcription Start Site (TSS). The resulting output is a unified vector consisting of continuous values that indicate the bulk RNA levels of a gene across 218 distinct tissue types.
@@ 3.5 Codebase Structure
Dataset HyenaDNA($\uparrow$) DNABERT($\uparrow$) DNABERT2($\uparrow$) GENA-LM($\uparrow$) NT($\uparrow$) Caduceus($\uparrow$) CNN($\uparrow$) Mouse Enhancers 0.7934 0.8099 0.8182 0.8297 0.8512 0.8163 0.7008 Coding vs Intergenomic 0.9097 0.9364 0.9358 0.9324 0.9576 0.9372 0.8844 Human vs Worm 0.9624 0.9584 0.9739 0.9698 0.9751 0.9557 0.9408 Human Enhancers Cohn 0.7296 0.7023 0.7587 0.7563 0.7612 0.7376 0.7080 Human Enhancers Ensembl 0.9033 0.8919 0.9075 0.9107 0.9244 0.8448 0.7637 Human Ensembl Regulatory 0.8462 0.938 0.8832 0.8810 0.9403 0.7367 0.8616 Human Nontata Promoters 0.9445 0.8713 0.9524 0.9660 0.9295 0.8885 0.8564 Human OCR Ensembl 0.7914 0.7496 0.7582 0.7898 0.8042 0.8176 0.6947 Human Core Promoter Detection 0.8440 0.8491 0.8257 0.8140 0.8541 0.8505 0.8003 Human Transcription Factor Prediction 0.6976 0.7840 0.8218 0.8240 0.8262 0.6928 0.6672 Human Promoter Detection 0.7295 0.8393 0.8993 0.9001 0.9390 0.7322 0.6875 Human Splice Site Detection 0.5660 0.8721 0.8813 0.9178 0.9481 0.5674 0.5666 Mouse Transcription Factor Prediction 0.6535 0.7393 0.8269 0.8265 0.8502 0.6519 0.6081 Yeast Epigenetic Marks Prediction 0.6301 0.7203 0.8022 0.7829 0.7845 0.6378 0.6071 Virus Covid Variant Classification 0.3770 0.5990 0.7195 0.7033 0.6939 0.3794 0.1974
@@ 4.1 Short Range Tasks
Convolution-based models consistently show lower performance compared to attention-based models, particularly on challenging tasks. In binary classification scenarios, Hyena-DNA demonstrates a notable decrease in accuracy, exhibiting a reduction of 0.053 compared to the attention-based models. This discrepancy becomes even more pronounced in multi-class classification tasks, where the accuracy gap widens to 0.239. Similarly, another model, Caduceus, shows a comparable pattern, with a 0.053 accuracy gap in binary classification tasks and a substantially larger margin of 0.274 in multi-class classification assignments. When comparing these models to the performance of pre-trained models, the discrepancy becomes even more striking. The CNN model, in particular, achieved lower accuracy, with the accuracy gap reaching up to 0.387. These findings highlight the challenges faced by convolution-based models like Hyena-DNA and Caduceus, especially when compared to attention-based and pre-trained models.
@@ 4.3 Effect of Length
We conducted an extensive analysis to evaluate the impact of sequence length on model performance in long-range genomic tasks. Specifically, we used input sequences of varying lengths—512, 1000, 2000, and 3000 base pairs (bp)—to assess the performance of four models: Hyena-DNA, Nucleotide Transformer, DNABERT-2, and GENA-LM in both species and promoter prediction tasks. Additionally, to explore the potential of convolution-based models, we tested a significantly longer input sequence of 30,000 bp with Hyena-DNA, focusing exclusively on the species prediction task. The results of these analyses are depicted in the top right and left sections of Figure 2. From the data, it is evident that increasing the sequence length consistently enhances performance across all the models tested. This trend is particularly pronounced with Hyena-DNA, which, despite trailing behind attention-based models at shorter context lengths, exhibits superior performance with longer contexts. This improvement underscores the advantages of using extended context lengths in genomic sequence analysis. However, this benefit is not without its challenges. In tasks like promoter prediction, where input length is inherently capped at 3000 bp, Hyena-DNA’s reliance on longer sequences becomes a limiting factor. This limitation presents a significant area for future research, aiming to optimize model performance within these constraints and potentially develop novel approaches to leverage longer sequences more effectively within the confines of specific genomic datasets.
@@ 4.4 Gene Clustering
In this section, we examine the fine-tuned embedding models HyenaDNA, DNABERT2, GENA-LM, Caduceus, and Nucleotide Transformer. These models are utilized to encode gene sequences from various species. To visualize the embeddings, we extract the representations from the final hidden layer of each model and apply t-distributed Stochastic Neighbor Embedding (t-SNE) [29]. The visualization, presented in Figure 3, reveals clear clusters that offer both visual and quantitative insights. For instance, the Nucleotide Transformer (NT), which demonstrates the highest accuracy among the models, shows well-separated embeddings for distinct species, indicating effective differentiation. In contrast, HyenaDNA, which has the lowest accuracy, displays less differentiation among the embeddings of different species, suggesting that its representations are less distinct. This visualization underscores the varying capabilities of distinguishing between gene sequences from different species, with NT excelling in accuracy and clarity of separation, while HyenaDNA struggles in comparison. From the results, it is clear that the k-mer based approaches have a more significant advantage.
@@ 5 Conclusion and Dicussion
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
  • Did you include the license to the code and datasets? [Yes] See Section A.
  • Did you include the license to the code and datasets? [No] The code and the data are proprietary.
  • Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No] The used code and data are open-source and under the MIT license for research usage.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The used data has undergone ethical review.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Codebase Overview
In this section, we present a comprehensive overview of the codebase structure of GenBench. The codebase is organized into three abstracted layers, namely the core layer, algorithm layer, and user interface layer, arranged from the bottom to the top, as illustrated in Figure 5. Our codebase is under Apache-2.0 license, like HyenaDNA [7].
@@
Core Layer The core layer of GenBench includes key elements like data loaders for supported datasets, fundamental modules for supported models, and metrics for evaluation. Data loaders provide a standardized way of loading and preprocessing data. The modules contain essential unit implementations of supported models. Metrics offer a consistent method for evaluating results. This core layer sets the groundwork for upper layers to ensure adaptable usage.
@@ Appendix B Implementation Details
The table presented in Table 9 outlines the hyperparameters utilized in the various models supported across different datasets. Each model’s hyperparameters consist of layers, a width of the hidden dimension, parameter size, learning rate, embed dropout, residual dropout, optimizer, optimizer momentum, training epochs, batch size, LR scheduler, and reverse complement augmentation. It is important to note that sequence length is task-dependent and not directly related to the model.
B.1 Model Description